# Fast and reliable confidence intervals for a variance component or proportion

### v1

@@ Abstract
We show that confidence intervals for a variance component or proportion, with asymptotically correct uniform coverage probability, can be obtained by inverting certain test-statistics based on the score for the restricted likelihood. The results apply in settings where the variance or proportion is near or at the boundary of the parameter set. Simulations indicate the proposed test-statistics are approximately pivotal and lead to confidence intervals with near-nominal coverage even in small samples. We illustrate our methods’ application in spatially-resolved transcriptomics where we compute approximately 15,000 confidence intervals, used for gene ranking, in less than 4 minutes. In the settings we consider, the proposed method is between two and 28,000 times faster than popular alternatives, depending on how many confidence intervals are computed.
@@ 1 Introduction
Variance component models are widely used in applied statistics. Often, researchers are interested in quantifying how much of the variability in the data is due to a specific source, possibly after controlling for, or conditioning on, other sources of variability. In modern genetics and genomics, for example, it is common to assume a vector $Y\in\mathbb{R}^{n}$ of responses for $n$ sample units satisfies
@@
with $X\in\mathbb{R}^{n\times p}$ a design matrix, $K\in\mathbb{R}^{n\times n}$ a known positive semi-definite matrix, and parameter $\theta=(\beta,\sigma_{g}^{2},\sigma_{e}^{2})\in\mathbb{R}^{p}\times[0,\infty)\times(0,\infty)$ (Purcell,, 2002; Kang et al.,, 2010; Zhang et al.,, 2010; Lippert et al.,, 2011; Yang et al.,, 2011; Zhou and Stephens,, 2012; Zhou et al.,, 2013; Loh et al.,, 2015; Weissbrod et al.,, 2016; Moore et al.,, 2019). In statistical genetics, when elements of $Y$ are phenotypes for $n$ individuals, $K$ can be a genetic relatedness matrix quantifying similarities between individuals’ common SNP genotypes (Jiang et al.,, 2022). Then, the quantity of inferential interest is the heritability (Visscher et al.,, 2008)
$h^{2}=\frac{\sigma_{g}^{2}}{\sigma_{g}^{2}+\sigma_{e}^{2}}\in[0,1).$ (2)
In spatially resolved transcriptomics (Marx,, 2021), when elements of $Y$ are a gene’s expression at $n$ distinct two- or three-dimensional spatial positions in a tissue sample and $K$ is a matrix encoding distances between positions, $h^{2}$ is known as the fraction of spatial variance (Svensson et al.,, 2018; Kats et al.,, 2021; Weber et al.,, 2022; Zeng et al.,, 2022).
@@
Though (1) may seem a simple model, inference on $h^{2}$ can be complicated, especially when it is nearly, or equal to, zero or one, as those are boundary points of the parameter set. It is well-known that common test-statistics, such as Wald and likelihood ratio statistics, have non-standard distributions on boundary points. The distributions can often be obtained by simulation (e.g., Crainiceanu and Ruppert,, 2004; Schweiger et al.,, 2016; Guédon et al.,, 2023) or approximated, either using asymptotic theory (e.g., Self and Liang,, 1987; Geyer,, 1994; Stram and Lee,, 1994, 1995; Baey et al.,, 2019), or by other means (e.g. Wood,, 2013). Thus, it is often possible to construct tests with (asymptotically) correct pointwise size, and the likelihood ratio test has recently been suggested for use in related settings (Battey and McCullagh,, 2023). However, creating a reliable confidence region by inverting such a test-statistic is more complicated, for at least two reasons.
@@
First, one in general has to compute the statistic for every relevant $h^{2}$. Thus, methods which are designed specifically for testing $h^{2}=0$ are inapplicable, and simulation-based methods can be prohibitively time-consuming. Secondly, even when inverting a test with asymptotically correct pointwise size is computationally feasible, the resulting confidence region can have poor coverage properties in finite samples. Specifically, the actual coverage probability often depends substantially on how close the true parameter is to the boundary. To illustrate the issues, Fig. 1 shows coverage probabilities for different values of $h^{2}$ for Wald, likelihood ratio, and proposed confidence intervals defined in Section  2. For interior points of the parameter set, the confidence intervals are obtained by inverting test-statistics using quantiles of the chi-square distribution with one degree of freedom, as suggested by classical theory. Notably, the coverage probabilities for Wald and likelihood ratio intervals are substantially different from the nominal level when $h^{2}$ is near zero or one. The reason is that, when the true parameter is close to the boundary, the distributions of the test-statistics are close to what they are on the boundary. Thus, even when the parameter is in the interior, classical asymptotic theory for interior points provides a poor approximation. Removing the boundary points from the parameter set or using a larger sample size is not sufficient to address the issue, as Fig. 1 illustrates; loosely speaking, for every sample size, there are interior points close enough to the boundary to cause problems. More formally, test-statistics with different asymptotic distributions on boundary and interior points lead to asymptotically incorrect uniform coverage probability (Ekvall and Bottai,, 2022, Lemma 2.5).
@@
Figure 1: Monte Carlo estimates of coverage probabilities for $h^{2}$ in (1) with $\sigma^{2}_{g}+\sigma_{e}^{2}=1$, $\beta=0$, $X\in\mathbb{R}^{n\times 5}$ a matrix of independent standard normal entries, and $K_{ij}=0.95^{|i-j|}$. The solid horizontal line indicates the nominal level, 0.95. The line widths provide 95% confidence bands for the coverage probability based on $10^{4}$ trials.
@@
For reasons to become apparent, we focus on test-statistics based on the score and information of the restricted likelihood. By doing so we are able to construct a $(1-\alpha)$ confidence region $\mathcal{R}_{n}(\alpha)$ that satisfies, for any compact $C\subseteq[0,1]$ and as $n\to\infty$,
@@
That is, the proposed confidence region, which is typically an interval, has asymptotically correct uniform coverage probability on compact sets. This property leads to near-nominal coverage of every practically relevant $h^{2}$, including $h^{2}$ near or equal to zero, as is illustrated in Fig. 1. We emphasize that when focusing on uniform inference in the presence of boundary points, conclusions about score test-statistics in general do not carry over to Wald and likelihood ratio test-statistics. Indeed, our results show, in a sense to be made precise, that certain score-statistics are asymptotically pivotal, but Wald and likelihood ratio test-statistics are not. The score statistics can therefore be inverted using the same quantiles for every point in the parameter set, leading to a confidence interval that is easy to implement and orders of magnitude faster to compute than competing ones. In a typical example we consider, our confidence intervals take less than a second to compute, whereas the competitors often take 500 seconds, and sometimes much longer.
Previously, the score standardized by expected information has been shown to give confidence regions with asymptotically correct uniform coverage probability in settings with singular Fisher information (Bottai,, 2003; Ekvall and Bottai,, 2022). We build on ideas from that literature here, but the setting, and hence the arguments required for the main results, are quite different. In particular, here, the information matrix is positive definite under weak conditions (Theorem 1). There are also two nuisance parameters, $\beta$ and $\sigma^{2}=\sigma_{g}^{2}+\sigma_{e}^{2}$. While dealing with $\beta$ is relatively straightforward since $\beta$ and $(h^{2},\sigma^{2})$ are orthogonal in the Fisher information sense, our results are, to the best of our knowledge, the first of their kind with non-orthogonal nuisance parameters. In contrast to classical asymptotic theory for inference in mixed models, we also let the number of predictors grow with the sample size. This is an important development as the restricted likelihood is typically used precisely when the number of predictors is smaller than the number of observations, but large enough to induce substantial bias in the inference on $h^{2}$.
@@ 2 Model
$l(\beta,\sigma_{g}^{2},\sigma_{e}^{2})=l(\beta,\sigma_{g}^{2},\sigma_{e}^{2};Y,X,K)=-\frac{1}{2}\left\{\log|\Sigma|+(Y-X\beta)^{\mathrm{\scriptscriptstyle T}}\Sigma^{-1}(Y-\beta)\right\},$ (4)
@@
$l(\beta,h^{2},\sigma^{2})=-\frac{1}{2}\sum_{i=1}^{n}\left\{\log(\sigma^{2})+\log(\lambda_{i}h^{2}+1-h^{2})+\frac{(Y_{i}-X_{i}^{\mathsf{T}}\beta)^{2}}{\sigma^{2}(\lambda_{i}h^{2}+1-h^{2})}\right\},$ (5)
where $X_{i}^{\mathsf{T}}$ is the $i$th row of $X$. The domain of $l$ is the parameter set $\mathbb{R}^{p}\times[0,1)\times(0,\infty)$. The corresponding restricted log-likelihood, with domain $[0,1)\times(0,\infty)$, is
@@ 2.1 Test-statistics and confidence regions
There are many potential test-statistics for inference based on $l(\beta,h^{2},\sigma^{2})$ and $l_{R}(h^{2},\sigma^{2})$. However, most common ones, like Wald and likelihood ratio statistics, do not give confidence regions satisfying (3) in boundary settings in general. Therefore, we focus on score statistics standardized to have mean zero and diagonal covariance matrix. Such test-statistics have previously been shown to lead to (3) in settings where, unlike here, the Fisher information is singular (Ekvall and Bottai,, 2022). Because our goal is to construct confidence regions for the covariance parameters, effectively treating $\beta$ as a nuisance parameter, we primarily work with the restricted likelihood.
Let $V_{1}=V_{1}(h^{2},\sigma^{2})=\partial\Sigma/\partial h^{2}=\sigma^{2}O(\Lambda-I_{n})O^{\mathrm{\scriptscriptstyle T}}$ and $V_{2}=V_{2}(h^{2},\sigma^{2})=\partial\Sigma/\partial\sigma^{2}=O(h^{2}\Lambda+(1-h^{2})I_{n})O^{\mathrm{\scriptscriptstyle T}}$. Similarly, let $U_{1}(h^{2},\sigma^{2})=\partial l_{R}(h^{2},\sigma^{2})/\partial h^{2}$ and $U_{2}(h^{2},\sigma^{2}_{e})=\partial l_{R}(h^{2},\sigma^{2})/\partial\sigma^{2}$, so the restricted score is $U(h^{2},\sigma^{2})=[U_{1}(h^{2},\sigma^{2}),U_{2}(h^{2},\sigma^{2})]^{\mathrm{\scriptscriptstyle T}}$. Routine calculations show, for $j\in\{1,2\}$,
@@
Let $\mathcal{I}(h^{2},\sigma^{2})$ denote the restricted information matrix, that is, the covariance matrix of $U(h^{2},\sigma^{2})$ when $h^{2}$ and $\sigma^{2}$ are the true parameters. We next give an expression for $\mathcal{I}(h^{2},\sigma^{2})$ for easy reference. The proof, along with those of other formally stated results, is in Appendix A.
Let $P=\Sigma^{-1/2}X(X^{\mathrm{\scriptscriptstyle T}}\Sigma^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\Sigma^{-1/2}$, $Q=I_{n}-P$, $D=\{h^{2}\Lambda+(1-h^{2})I_{n}\}^{-1}(\Lambda-I_{n})$, and $H=ODO^{\mathsf{T}}$.
@@
For inference on $(h^{2},\sigma^{2})$ jointly, we consider the (quadratic) restricted score standardized by the restricted Fisher information:
@@ 3 Asymptotic distributions
We assume $K_{n}=\Lambda_{n}=\operatorname{diag}(\lambda_{n1},\dots,\lambda_{nn})$ in this section without loss of generality, and hence $\Sigma_{n}=\sigma^{2}_{n}(I_{n}+\Lambda_{n})$ and $H_{n}=D_{n}$ are diagonal. Here, a subscript $n$ has been added to $\sigma^{2}$ and other quantities since we will consider a sequence $\{(\beta_{n},h^{2}_{n},\sigma^{2}_{n})\}$ of parameters indexed by the sample size, and a corresponding sequence of models. Similarly, the number of predictors $p=p_{n}$ can change with $n$. By considering a sequence of (true) parameters, we can obtain the asymptotic distribution of $T^{R}_{n}(h^{2}_{n},\sigma^{2}_{n})$ and $T_{n}^{R}(h_{n}^{2})$ under that sequence. This is useful because (3) holds for $\mathcal{R}_{n}(\alpha)$ in (10) if and only if the asymptotic distribution of $T_{n}^{R}(h_{n}^{2})$ is the same under any convergent sequence of parameters (Ekvall and Bottai,, 2022, Lemma 2.5), and similarly for $\mathcal{R}^{2}_{n}(\alpha)$. This result essentially follows from the equivalence of continuous convergence and uniform convergence on compact sets, applied to the sequence of functions defined by $\theta\mapsto\mathrm{P}_{\theta}\{T_{n}^{R}(h^{2})\leq q_{1-\alpha,1}\}$. More generally, examining the distribution under a sequence of parameters tending to a point, can lead to a better understanding of the behavior of the test-statistic near that point. Thus, for our purposes, sequences tending to boundary points are of particular interest. The next result is our first along these lines.
@@
Condition (ii) says the sum of the $p$ greatest eigenvalues of $D_{n}^{2}$ should be negligible compared to the sum of all its eigenvalues. An eigenvalue, or diagonal element, of $D_{n}^{2}$ will be relatively large if either (a) the corresponding element of $\Lambda_{n}$ is relatively close to zero and $h^{2}$ is close to one, or (b) the corresponding element of $\Lambda_{n}$ is relatively large. Thus, condition (ii) suggests the test-statistic may not behave well if a few elements of $O^{\mathrm{\scriptscriptstyle T}}Y$ have much larger variance than all other ones, or if some elements have very small variance and most if it is attributable to $K$.

### v2

@@ Abstract
We show that confidence intervals in a variance component model, with asymptotically correct uniform coverage probability, can be obtained by inverting certain test-statistics based on the score for the restricted likelihood. The results apply in settings where the variance is near or at the boundary of the parameter set. Simulations indicate the proposed test-statistics are approximately pivotal and lead to confidence intervals with near-nominal coverage even in small samples. We illustrate our methods’ application in spatially-resolved transcriptomics where we compute approximately 15,000 confidence intervals, used for gene ranking, in less than 4 minutes. In the settings we consider, the proposed method is between two and 28,000 times faster than popular alternatives, depending on how many confidence intervals are computed.
@@ 1 Introduction
Variance component models are widely used in applied statistics. Often, researchers are interested in quantifying how much of the variability in the data is due to a specific source, possibly after controlling for, or conditioning on, other sources of variability. It is common to assume that for non-stochastic $X\in\mathbb{R}^{n\times p}$ and $Z\in\mathbb{R}^{n\times q}$, a vector of responses $Y\in\mathbb{R}^{n}$ satisfies
$Y=X\beta+ZU+E,$ (1)
where $U\sim\mathrm{N}_{q}(0,\sigma_{g}I_{q})$ and, independently, $E\sim\mathrm{N}_{n}(0,\sigma^{2}_{e}I_{n})$. Both $U$ and $E$ are unobservable, and the parameter is $\theta=(\beta,\sigma^{2}_{g},\sigma^{2}_{e})\in\mathbb{R}^{p}\times[0,\infty)\times(0,\infty)$. Equation (1) is a linear mixed model with one random effect, or several independent random effects with the same variance.
In modern genetics and genomics, the model is often written directly for the marginal distribution of $Y$:
@@
where $K=ZZ^{\mathrm{\scriptscriptstyle T}}$; see for example Purcell, (2002); Kang et al., (2010); Zhang et al., (2010); Lippert et al., (2011); Yang et al., (2011); Zhou and Stephens, (2012); Zhou et al., (2013); Loh et al., (2015); Weissbrod et al., (2016); Moore et al., (2019). Equations (1) and (2) give equivalent models since, for any positive semi-definite $K\in\mathbb{R}^{n\times n}$ of rank $q\leq n$, there exists a $Z\in\mathbb{R}^{n\times q}$ such that $K=ZZ^{\mathrm{\scriptscriptstyle T}}$. In statistical genetics, when elements of $Y$ are phenotypes for $n$ individuals, $K$ can be a genetic relatedness matrix quantifying similarities between individuals’ common SNP genotypes (Jiang et al.,, 2022). Then, the quantity of inferential interest is the heritability (Visscher et al.,, 2008)
$h^{2}=\frac{\sigma_{g}^{2}}{\sigma_{g}^{2}+\sigma_{e}^{2}}\in[0,1).$
In spatially resolved transcriptomics (Marx,, 2021), when elements of $Y$ are a gene’s expression at $n$ distinct two- or three-dimensional spatial positions in a tissue sample and $K$ is a matrix encoding distances between positions, $h^{2}$ is known as the fraction of spatial variance (Svensson et al.,, 2018; Kats et al.,, 2021; Weber et al.,, 2022; Zeng et al.,, 2022). The parameter $h^{2}$ is related to $\tau=\sigma^{2}_{g}/\sigma^{2}_{e}$, which has been the focus of previous research on confidence intervals under (1) (e.g., Crainiceanu and Ruppert,, 2004), as $h^{2}=\tau/(1+\tau)$. Because the naming of parameters is inconsistent in the literature, we refer to all of $\sigma_{g}^{2}$, $\tau$, and $h^{2}$ as variance components for simplicity. To be concrete, we will focus on $h^{2}$ as that is common in the motivating applications, but similar results often apply in different parameterizations. A further discussion of how our results could be extended to other parameterizations and models, including ones with more variance components or a parameterized $K$, is in the Supplementary Material.
@@
Though (2) may seem a simple model, inference on $h^{2}$ can be complicated, especially when it is nearly, or equal to, zero or one, as those are boundary points of the parameter set. It is well-known that common test-statistics, such as Wald and likelihood ratio statistics, have non-standard distributions on boundary points. The distributions can often be obtained by simulation (e.g., Crainiceanu and Ruppert,, 2004; Schweiger et al.,, 2016; Guédon et al.,, 2024) or approximated, either using asymptotic theory (e.g., Self and Liang,, 1987; Geyer,, 1994; Stram and Lee,, 1994, 1995; Baey et al.,, 2019), or by other means (e.g. Wood,, 2013). Thus, it is often possible to construct tests with (asymptotically) correct pointwise size, and the likelihood ratio test has recently been suggested for use in related settings (Battey and McCullagh,, 2023). However, creating a reliable confidence interval by inverting such a test-statistic is more complicated, for at least two reasons.
@@
First, one in general has to compute the statistic for every relevant $h^{2}$. Thus, methods designed specifically for testing $h^{2}=0$ are inapplicable, and simulation-based methods can be prohibitively time-consuming. Secondly, even when inverting a test with asymptotically correct pointwise size is computationally feasible, the resulting confidence region can have poor coverage properties. Specifically, the actual coverage probability often depends substantially on how close the true parameter is to the boundary. More formally, test-statistics with different asymptotic distributions on boundary and interior points lead to asymptotically incorrect uniform coverage probability (Ekvall and Bottai,, 2022, Lemma 2.5).
[…]
@@
Figure 1: Monte Carlo estimates of coverage probabilities for $h^{2}$ in (2) with $\sigma^{2}_{g}+\sigma_{e}^{2}=1$, $\beta=0$, $X\in\mathbb{R}^{n\times 5}$ a matrix of independent standard normal entries, and $K_{ij}=0.95^{|i-j|}$. The solid horizontal line indicates the nominal level. The line widths provide 95% confidence bands for the coverage probability based on $10^{4}$ trials.
@@
For reasons to become apparent, we focus on test-statistics based on the score and information of the restricted likelihood. By doing so we are able to construct a $(1-\alpha)$ confidence region $\mathcal{R}_{n}(\alpha)$ that satisfies, for any compact subset of the parameter set, say $C$, and as $n\to\infty$,
@@
That is, the proposed confidence region, which is typically an interval, has asymptotically correct uniform coverage probability on compact sets. This property leads to near-nominal coverage of every practically relevant $h^{2}$, including $h^{2}$ near or equal to zero, as is illustrated in Fig. 1. We emphasize that when focusing on uniform inference in the presence of boundary points, conclusions about score test-statistics in general do not carry over to Wald and likelihood ratio test-statistics. Indeed, our results show, in a sense to be made precise, that certain score statistics are asymptotically pivotal, but Wald and likelihood ratio test-statistics are not. The score statistics can therefore be inverted using the same quantiles for every point in the parameter set, leading to a confidence interval that is easy to implement and orders of magnitude faster to compute than competing ones. In a typical example we consider, our confidence intervals take less than a second to compute, whereas the competitors often take 500 seconds, and sometimes much longer.
Previously, the score standardized by expected information has been shown to give confidence regions with asymptotically correct uniform coverage probability in settings with singular Fisher information (Bottai,, 2003; Ekvall and Bottai,, 2022). We build on similar ideas here, but the setting, and hence the arguments required for the main results, are quite different. In particular, here, the information matrix is positive definite under weak conditions (Theorem 1). There are also two nuisance parameters, $\beta$ and $\sigma^{2}$. While dealing with $\beta$ is relatively straightforward since $\beta$ and $(h^{2},\sigma^{2})$ are orthogonal in the Fisher information sense, our results are, to the best of our knowledge, the first of their kind with non-orthogonal nuisance parameters. For instance, Ekvall and Bottai, (2022) assume $\sigma^{2}_{e}>0$ is known. Moreover, here, the number of predictors can grow with the sample size. This is an important development as the restricted likelihood is typically used precisely when the number of predictors is smaller than the number of observations, but large enough to induce substantial bias in the inference on $h^{2}$.
@@ 2.1 Likelihood functions
$l(\beta,\sigma_{g}^{2},\sigma_{e}^{2})=l(\beta,\sigma_{g}^{2},\sigma_{e}^{2};Y,X,K)=-\frac{1}{2}\left\{\log|\Sigma|+(Y-X\beta)^{\mathrm{\scriptscriptstyle T}}\Sigma^{-1}(Y-X\beta)\right\},$ (4)
@@
$l(\beta,h^{2},\sigma^{2})=-\frac{1}{2}\sum_{i=1}^{n}\left\{\log(\sigma^{2})+\log(\lambda_{i}h^{2}+1-h^{2})+\frac{(Y_{i}-X_{i}^{\mathrm{\scriptscriptstyle T}}\beta)^{2}}{\sigma^{2}(\lambda_{i}h^{2}+1-h^{2})}\right\},$ (5)
where $X_{i}^{\mathrm{\scriptscriptstyle T}}$ is the $i$th row of $X$. The domain of $l$ is the parameter set $\mathbb{R}^{p}\times[0,1)\times(0,\infty)$. The corresponding restricted log-likelihood, with domain $[0,1)\times(0,\infty)$, is
@@ 2.2 Test-statistics and confidence regions
There are many potential test-statistics for inference based on $l(\beta,h^{2},\sigma^{2})$ and $l_{R}(h^{2},\sigma^{2})$. However, most common ones, like Wald and likelihood ratio statistics, do not give confidence regions satisfying (3) in boundary settings in general. Therefore, we focus on score statistics standardized to have mean zero and diagonal covariance matrix. Such test-statistics have previously been shown to lead to results like (3) in settings where, unlike here, the Fisher information is singular (Ekvall and Bottai,, 2022). Because our goal is to construct confidence regions for the covariance parameters, effectively treating $\beta$ as a nuisance parameter, we primarily work with the restricted likelihood.
Let $V_{1}=V_{1}(h^{2},\sigma^{2})=\partial\Sigma/\partial h^{2}=\sigma^{2}O(\Lambda-I_{n})O^{\mathrm{\scriptscriptstyle T}}$ and $V_{2}=V_{2}(h^{2},\sigma^{2})=\partial\Sigma/\partial\sigma^{2}=O(h^{2}\Lambda+(1-h^{2})I_{n})O^{\mathrm{\scriptscriptstyle T}}$. Similarly, let $U_{1}(h^{2},\sigma^{2})=\partial l_{R}(h^{2},\sigma^{2})/\partial h^{2}$ and $U_{2}(h^{2},\sigma^{2})=\partial l_{R}(h^{2},\sigma^{2})/\partial\sigma^{2}$, so the restricted score is $U(h^{2},\sigma^{2})=[U_{1}(h^{2},\sigma^{2}),U_{2}(h^{2},\sigma^{2})]^{\mathrm{\scriptscriptstyle T}}$. Routine calculations show, for $j\in\{1,2\}$,
@@
Let $\mathcal{I}(h^{2},\sigma^{2})$ denote the restricted information matrix, that is, the covariance matrix of $U(h^{2},\sigma^{2})$ when $h^{2}$ and $\sigma^{2}$ are the true parameters. We next give an expression for $\mathcal{I}(h^{2},\sigma^{2})$ for easy reference. The proof, along with those of other formally stated results, is in the Supplementary Material.
Let $P=\Sigma^{-1/2}X(X^{\mathrm{\scriptscriptstyle T}}\Sigma^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\Sigma^{-1/2}$, $Q=I_{n}-P$, $D=\{h^{2}\Lambda+(1-h^{2})I_{n}\}^{-1}(\Lambda-I_{n})$, and $H=ODO^{\mathrm{\scriptscriptstyle T}}$.
@@
For inference on $(h^{2},\sigma^{2})$ jointly, we consider the (quadratic) restricted score standardized by the restricted information:
@@ 3 Asymptotic distributions
We assume $K_{n}=\Lambda_{n}=\operatorname{diag}(\lambda_{n1},\dots,\lambda_{nn})$ in this section without loss of generality, and hence $\Sigma_{n}=\sigma^{2}_{n}\{(1-h_{n})I_{n}+h_{n}\Lambda_{n}\}$ and $H_{n}=D_{n}$ are diagonal. Here, a subscript $n$ has been added to $\sigma^{2}$ and other quantities since we will consider a sequence $\{(\beta_{n},h^{2}_{n},\sigma^{2}_{n})\}$ of parameters indexed by the sample size, and a corresponding sequence of models. Similarly, the number of predictors $p=p_{n}$ can change with $n$. By considering a sequence of (true) parameters, we can obtain the asymptotic distribution of $T^{R}_{n}(h^{2}_{n},\sigma^{2}_{n})$ and $T_{n}^{R}(h_{n}^{2})$ under that sequence. This is useful because (3) holds for $\mathcal{R}_{n}(\alpha)$ in (10) if and only if the asymptotic distribution of $T_{n}^{R}(h_{n}^{2})$ is the same under any convergent sequence of parameters (Ekvall and Bottai,, 2022, Lemma 2.5), and similarly for $\mathcal{R}^{2}_{n}(\alpha)$. This result essentially follows from the equivalence of continuous convergence and uniform convergence on compact sets, applied to the sequence of functions defined by $\theta\mapsto\mathrm{P}_{\theta}\{T_{n}^{R}(h^{2})\leq q_{1-\alpha,1}\}$. More generally, examining the distribution under a sequence of parameters tending to a point, can lead to a better understanding of the behavior of the test-statistic near that point. Thus, for our purposes, sequences tending to boundary points are of particular interest. The next result is our first along these lines.
@@
Condition (ii) says the sum of the $p$ greatest eigenvalues of $D_{n}^{2}$ should be negligible compared to the sum of all its eigenvalues. An eigenvalue, or diagonal element, of $D_{n}^{2}$ will be relatively large if either (a) the corresponding element of $\Lambda_{n}$ is relatively close to zero and $h^{2}$ is close to one, or (b) the corresponding element of $\Lambda_{n}$ is relatively large. Thus, condition (ii) suggests the test-statistic may not behave well if a few elements of $O^{\mathrm{\scriptscriptstyle T}}Y$ have much larger variance than all other ones, or if some elements have very small variance and most of it is attributable to $K$.